1) (2 pts) a) Jan. 1, 2014, the DJIA was at 15,300. On Jan. 1, 2019, the DJIA is at 23,800. What was the geometric average annual rate of return over the last 5 years? 13.5% Stock BB currently sells for $49 per share and is expected to pay a $1.00 dividend over nths and sell for $53 in 9 months. What is the expected periodic return over the 2) 2 pts) t 9 months ? What is the expected nominal rate of return (Annual Percentage Rate) for Stock nex BB? Expected Periodic Rate of Return% Expected Nominal Rate of Return (APR)- % - % (2 pts) Stock CC currently sells for $42 a share and is expected to pay $1.00 dividend one year from now and $1.80 dividend two years from now and $2.10 dividend three years from now. Stock CC is expected to sell for $69 per share in three years. What is the expected rate of return for Stock CC? 3) Expected rate of return for Stock CC 8.9
Answer #1

Answer:

Question 1)

1st January 2014 DJIA P=15300

1st January 2019 DJIA A=23800

n=5 years

Let Geometric average annual rate= r

So A=P*(1+r/100)^n

23800=15300*(1+r/100)^5

(1+r/100)^5=1.56

r=9.30%